# The systemic recoil velocity distribution and the scale height of isolated millisecond pulsar systems: Implications on neutron star retention fractions in star clusters

### v1

@@ 1 Introduction
Up till now, a well defined MSP $v_{\mathrm{sys}}$ distribution remains largely absent, mainly due to the aforementioned technical challenges and relatively large astrometric uncertainties. This absence hinders the studies of MSP formation channels, and results in weak constraints on the NS retention fractions in star clusters. Recent years have seen a major increase in the number of well astrometrically determined MSPs (e.g. Ding et al., 2023; Shamohammadi et al., 2024), which promises a refined characterization of the MSP $v_{\mathrm{sys}}$ distribution. In this paper, we determine the MSP $v_{\mathrm{sys}}$ distribution with the observed $v_{\mathrm{l}}$ calculated for 65 well astrometrically determined MSPs, then discuss its implications for the NS retention fractions in star clusters. Throughout this paper, uncertainties are stated at 68% confidence.
The establishment of the observed $\bm{v}_{\perp}$ distribution requires precise determination of both proper motions and distances for the MSPs, especially as MSP $v_{\perp}$ are much lower than those of normal pulsars (Hobbs et al., 2005; Gonzalez et al., 2011). Applying low-precision astrometric results to MSP kinematics studies would not only shift the average $v_{\perp}$ to the higher end, but also reduce the ability to resolve the potential multiple modes in the $\bm{v}_{\perp}$ distribution. Compared to other methods, the determination of trigonometric parallaxes through VLBI astrometry (e.g. Ding et al., 2023), pulsar timing (e.g. Perera et al., 2019; Shamohammadi et al., 2024), or Gaia astrometry (e.g. Jennings et al., 2018; Moran et al., 2023), offers the most precise and reliable distances for MSPs. Recent releases of the astrometric results by Ding et al. (2023) and Shamohammadi et al. (2024) have significantly enriched the sample of well astrometrically determined MSPs, which offers a rare opportunity to make statistically significant investigation of the MSP kinematics based purely on well astrometrically determined MSPs.
We searched for viable MSPs within the PSRCAT Catalogue [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/ (Manchester et al., 2005). Due to relatively frequent dynamic interactions inside globular clusters, the $\bm{v}_{\perp}$ of globular cluster MSPs are likely not reflective of their $v_{\mathrm{sys}}$. With globular cluster MSPs excluded from the study, we reached 65 isolated MSP systems (sometimes referred to as field MSP systems) having significant ($>3\,\sigma$) proper motion and parallax determinations (see Table 2). The details of the sample selection are described in Appendix A.
Following the methods detailed in Section 6 of Ding et al. (2023), we derived the distances $d$ and the $\bm{v}_{\perp}$ of the 65 MSPs from their parallaxes and proper motions. From the $d$, we calculated the Galactic heights $z\equiv d\sin{b}$ for the 65 MSPs, where $b$ refer to Galactic latitudes. The derived $d$, $\bm{v}_{\perp}$ and $z$ are summarized in Table 2, where the columns related to $\bm{v}_{\perp}$ include the $\bm{v}_{\perp}$ magnitudes $v_{\perp}$, the Galactic-longitude component $v_{\mathrm{l}}$ and the Galactic-latitude component $v_{\mathrm{b}}$ of $\bm{v}_{\perp}$. On top of $v_{\mathrm{l}}$ and $v_{\mathrm{b}}$, $z$ serves as an additional indicator of MSP kinematics, as the average level of $|z|$ expected to grow with higher $v_{\mathrm{sys}}$. Therefore, this work includes the studies of three samples of measurements: the $v_{\mathrm{l}}$, the $v_{\mathrm{b}}$, and the $z$ samples.
Based on the calculations of the Pearson coefficients, no linear correlation is found between any two of the three samples. Accordingly, the three samples are treated independently. We conducted Wilcoxon signed-rank test on each of the three samples of 65 measurements, and confirmed no indication of asymmetry about 0 in any sample. Accordingly, in the following analysis, we assumed that the samples of the observed $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$ are symmetric about 0.
To 1) visualize the three samples of measurements, and 2) investigate the underlying probability density function (PDF) for each sample, we assumed each measurement follows a split normal distribution, simulated 10,000 values per measurement, and concatenated the simulations into a chain of 650,000 values for each of $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$. Despite the simulated sample of 650,000 following essentially a multi-modal distribution, the visualization is expected to characterize where the measurements overlap. The resultant histograms are displayed in Figure 2. According to Table 2 and Figure 2, most of the 65 MSPs are within the scale height of the Galactic thick disk (i.e., $\sim 1$ kpc).
Figure 2: The upper-left, lower-left and lower-right panels display the histograms of Galactic heights $z$, the Galactic-longitude component $v_{\mathrm{l}}$ and the Galactic-latitude component $v_{\mathrm{b}}$ of transverse peculiar velocities, respectively. Each histogram is concatenated from 10000 simulations drawn from the assumed split normal distributions for the 65 measurements (of $z$, $v_{\mathrm{l}}$ or $v_{\mathrm{b}}$; see Table 2). The gray dashed curves show the normalized count smoothed with kernel density estimation (using scipy.stats.Gaussian_kde). The best-fit Cauchy, Laplace and three-component normal distributions are plotted in blue and green dashed curves. Overlaid are the dynamical population synthesis results (labeled as “evolved”) that best match the observed $z$, $v_{\mathrm{l}}$ and $v_{\mathrm{b}}$, based on the determined initial MSP scale height $\zeta_{0}=0.32$ kpc. The corresponding initial distributions are also provided except for $v_{\mathrm{l}}$, as the initial $v_{\mathrm{l}}$ distribution adopted in the dynamical population synthesis is identical to the best-fit three-component normal distribution (a linear combination of three normal distributions).
Following the method detailed in Section 7.2.3 of Ding et al. (2024b), we ran Hartigan’s dip tests (Hartigan & Hartigan, 1985) on the three samples (of 65 measurements), and found no indication of multi-modality in any of the samples. Should the $v_{\mathrm{l}}$ sample truly follow the multi-modal distribution used in the visualization, about 200 and 260 additional $v_{\mathrm{l}}$ measurements would be needed to rule out uni-modality of the $|v_{\mathrm{l}}|$ sample at 90% and 95% confidence, respectively, according to our further simulations.
@@ 3 The systemic recoil velocity distribution
We followed Faucher-Giguere & Kaspi (2006) to approach the $v_{\mathrm{sys}}$ distribution with the observed $v_{\mathrm{l}}$ sample (see Table 2), assuming that $v_{\mathrm{l}}$ stays statistically stable over the long ($\lesssim 10$ Gyr) journeys of MSPs. The validation of this “stable-$v_{\mathrm{l}}$” assumption is detailed in Section 4. We determined the $v_{\mathrm{sys}}$ distribution in mainly three steps.
@@
In the second step, we fit the parameter(s) of each candidate PDF using scipy.optimize.curve_fit (Virtanen et al., 2020), for the PDF to approximate the curve (see Figure 2) smoothed from the normalized histogram of the measurement-based simulated $v_{\mathrm{l}}$ sample of 650,000 (see Section 2), where the smoothing was conducted with kernel density estimation using scipy.stats.Gaussian_kde (Virtanen et al., 2020). After discarding candidates with poor fitting results, we identified two groups of prime PDF candidates (that well match the normalized $v_{\mathrm{l}}$ histogram), with one group being linear combinations of Cauchy distributions and the other group being linear combinations of the normal distributions. The prime candidate PDFs are described in detail in Appendix B.1.
@@
In the final step, we selected the PDF for the observed $v_{\mathrm{l}}$ sample by the method detailed in Appendix B.2. We conclude that the observed $v_{\mathrm{l}}$ can be well described by a Cauchy distribution with scale of 51.8 $\mathrm{km\leavevmode\nobreak\ s^{-1}}$ (i.e., Equation 8). Despite a three-component normal distribution being more consistent with the Cauchy distribution (see Figure 2), the more complex distribution has four more free parameters, therefore being less favorable. On the other hand, the three-component normal distribution, described by Equation 3 and Table 4, not only serves as a good approximation to the underlying PDF of the observed $v_{\mathrm{l}}$ sample, but also provides a rare analytical solution for the $\bm{v}_{\mathrm{sys}}$ magnitude $v_{\mathrm{sys}}$: when assuming the $\bm{v}_{\mathrm{sys}}$ distribution is isotropic for kinematically old pulsar systems, $v_{\mathrm{sys}}$ is expected to follow a three-component Maxwell distribution (i.e., a weighted sum of three Maxwell distributions) that share the same parameters with the best-fit three-component normal distribution (see Table 4). Namely, under the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption, $v_{\mathrm{sys}}$ should approximately follow the PDF
$\begin{split}p=&\sqrt{\frac{2}{\pi}}\Biggl{\{}0.498\cdot\frac{{v_{\mathrm{sys}}}^{2}}{45.5^{3}}\cdot\exp{\left[-\frac{1}{2}\left(\frac{v_{\mathrm{sys}}}{45.5}\right)^{2}\right]}\\ &\quad+0.447\cdot\frac{{v_{\mathrm{sys}}}^{2}}{110.3^{3}}\cdot\exp{\left[-\frac{1}{2}\left(\frac{v_{\mathrm{sys}}}{110.3}\right)^{2}\right]}\\ &\quad+0.055\cdot\frac{{v_{\mathrm{sys}}}^{2}}{448.8^{3}}\cdot\exp{\left[-\frac{1}{2}\left(\frac{v_{\mathrm{sys}}}{448.8}\right)^{2}\right]}\Biggl{\}}\,\,,\end{split}$ (1)
@@
where the unit of $p$ and $v_{\mathrm{sys}}$ is $\mathrm{s\leavevmode\nobreak\ km^{-1}}$ and $\mathrm{km\leavevmode\nobreak\ s^{-1}}$, respectively. In contrast, no analytical solution of the $v_{\mathrm{sys}}$ PDF can be solved from the best-fit Cauchy distribution of $v_{\mathrm{l}}$. Therefore, we adopt Equation 1 as the $v_{\mathrm{sys}}$ PDF in the following analysis, which is illustrated along with its cumulative distribution function (CDF) in Figure 4.
@@
Regarding the use of Equation 1, we like to clarify two points. Firstly, despite being a three-component Maxwell distribution, Equation 1 is still a uni-modal distribution by definition (see Figure 4). Secondly, though approximating the $v_{\mathrm{sys}}$ PDF by Equation 1, we do not suggest that isolated MSP systems have multiple formation channels. To further the investigation of formation channels of isolated MSP systems, population synthesis studies based on various MSP formation channels are needed to generate the theoretical $v_{\mathrm{sys}}$ PDF templates that can be compared with our observation-based counterpart.
@@ 4 Dynamical population synthesis
When deriving the $v_{\mathrm{sys}}$ PDF, we assumed $v_{\mathrm{l}}$ to be statistically stable over time (see Section 3). We tested this “stable-$v_{\mathrm{l}}$” assumption using DPS that simulates the dynamics of isolated MSP systems in the Galaxy. Specifically, we like to examine if the current (or evolved) $v_{\mathrm{l}}$ distribution stays roughly the same as the initial $v_{\mathrm{l}}$ distribution. We used Equation 1 as the MSP $v_{\mathrm{sys}}$ PDF in the DPS analysis. The MWPotential2014 model, calculated and compiled into the galpy package by Bovy (2015), was adopted as the Galactic potential used in the DPS. Apart from the $v_{\mathrm{sys}}$ PDF and the Galactic potential, two other ingredients are essential the DPS analysis — the MSP age distribution, and the initial spatial distribution of isolated MSP systems in the Galaxy. The former ingredient is described in Appendix C.1, while the latter is introduced as follows.
@@ 4.1 The initial Galactic spatial distribution
The initial Galactic distribution of isolated MSP systems consists of two parts — the birth radial distribution (i.e., the spatial distribution projected to the Galactic plane) and the birth Galactic height distribution (hereafter referred to as the $z_{\mathrm{i}}$ distribution). For the former part, we adopted the pulsar birth radial distribution by Xie et al. (2024), which assumes that Galactic isolated radio pulsars are born in the spiral arms.
On the other hand, the $z_{\mathrm{i}}$ distribution of isolated MSP systems is largely unknown. As we found that the observed $z$ sample (see Table 2) can be well described by a Laplace distribution with the scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc (see Figure 2 and Appendix C.2.1), we assumed that $z_{\mathrm{i}}$ of Galaxy-wide isolated MSP systems also follow a Laplace distribution with the scale height of $\zeta_{0}$. In DPS, we realized that both the evolved $v_{\mathrm{b}}$ and the evolved $z$ change with $\zeta_{0}$ (see Appendix C.2.2). Therefore, by matching the evolved $v_{\mathrm{b}}$ and the evolved $z$ to the observed counterparts, one can determine $\zeta_{0}$. The DPS procedure that determines the evolved states (including the evolved positions and the evolved velocities) is described in Section 4.2, while the details of the $\zeta_{0}$ determination are provided in Appendix C.2.2. We find that the DPS with $\zeta_{0}=0.32$ kpc renders evolved $v_{\mathrm{b}}$ and evolved $z$ samples that well match the observed counterparts (see Figure 10). Accordingly, we adopt $\zeta_{0}=0.32$ kpc for the Galaxy wide $z_{\mathrm{i}}$ distribution of isolated MSP systems.
@@ 4.2 The procedure of dynamical population synthesis
Subsequently, we performed orbit integrations with galpy to obtain the evolved positions and velocities after the span of each age in the sample of 13,005 MSP ages described in Appendix C.1. In this way, we compiled $10,000\times 13,005=130$ million evolved states (i.e., positions and velocities for different MSP ages) of simulated MSPs, from which $\sim 4$% evolved states outside the edge of the Galaxy (i.e., beyond 292 kpc from the Galactic centre, Deason et al., 2020) were discarded. To take into account the selection effect that astrometrically well determined pulsars are predominantly located within $\lesssim 8$ kpc, we removed evolved states that are beyond the 8 kpc distance. Additionally, given that the observed MSPs compiled in this work are $\lesssim 2.5$ kpc from the Galactic plane (see Figure 2), we discarded the evolved states with $|z_{\mathrm{e}}|>2.5$ kpc, where $z_{\mathrm{e}}$ refer to the evolved Galactic heights. The remaining evolved states are $\sim 37$ million in number, which amount to $\sim 30$% of the Galaxy-wide evolved states.
@@ 4.3 Results of dynamical population synthesis
Moreover, from the Galaxy-wide $z_{\mathrm{e}}$ sample (of $\sim 125$ million), we found that the best-fit scale height $\zeta_{1}$ for isolated MSP systems in the whole Galaxy is 0.66 kpc, which is 27% larger than the observed scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc estimated from relatively nearby MSPs. Interestingly, unlike the observed $z$ sample that is best described by a Laplace distribution (see Appendix C.2.1), the Galaxy-wide $z_{\mathrm{e}}$ sample is best characterized by the Cauchy distribution with 0.47 kpc scale (see Figure 12).

### v3

@@ 2.2 Modal analysis based on different parallax significance
$\varpi/\sigma_{\varpi}$ >3>3 >4>4 >5>5 >6>6 >7>7
$N_{\mathrm{sample}}$ 64 53 45 38 30
Hartigan’s dip tests on $v_{\mathrm{l}}$
p-value 0.91 0.89 0.81 0.77 0.92
$N_{0.9}$ $\approx 490$ $\approx 360$ $\approx 220$ $\approx 150$ $\approx 200$
$N_{0.95}$ $\approx 610$ $\approx 450$ $\approx 270$ $\approx 180$ $\approx 250$
Hartigan’s dip tests on $|v_{\mathrm{l}}|$
p-value 0.64 0.57 0.43 0.42 0.33
$N_{0.9}$ $\approx 240$ $\approx 160$ $\approx 110$ $\approx 90$ $\approx 60$
$N_{0.95}$ $\approx 300$ $\approx 200$ $\approx 130$ $\approx 110$ $\approx 70$
@@ 3 The systemic recoil velocity distribution
We followed Faucher-Giguere06 to approach the $v_{\mathrm{sys}}$ distribution with the observed $v_{\mathrm{l}}$ sample (see Table 2), assuming that $v_{\mathrm{l}}$ stays statistically stable over the long ($\lesssim 10$ Gyr) journeys of MSPs. The validation of this “stable-$v_{\mathrm{l}}$” assumption is detailed in Section 4. We determined the $v_{\mathrm{sys}}$ distribution in mainly three steps.
@@
In the second step, we fit the parameter(s) of each candidate PDF using scipy.optimize.curve_fit (Virtanen20), for the PDF to approximate the curve (see Figure 2) smoothed from the normalized histogram of the measurement-based simulated $v_{\mathrm{l}}$ sample of 640,000 (see Section 2), where the smoothing was conducted with kernel density estimation using scipy.stats.Gaussian_kde (Virtanen20). After discarding candidates with poor fitting results, we identified two groups of prime PDF candidates (that well match the normalized $v_{\mathrm{l}}$ histogram), with one group being linear combinations of Cauchy distributions and the other group being linear combinations of the normal distributions. The prime candidate PDFs are described in detail in Appendix B.1.
@@
In the final step, we selected the PDF for the observed $v_{\mathrm{l}}$ sample by the method detailed in Appendix B.2. We conclude that the observed $v_{\mathrm{l}}$ can be well described by a Cauchy distribution with scale of 51.2 $\mathrm{km\penalty 10000\ s^{-1}}$ (i.e., Equation 8). Despite a three-component normal distribution being more consistent with the Cauchy distribution (see Figure 2), the more complex distribution has four more free parameters, therefore being less favorable. On the other hand, the three-component normal distribution, described by Equation 3 and Table 8, not only serves as a good approximation to the underlying PDF of the observed $v_{\mathrm{l}}$ sample, but also provides a rare analytical solution for the $\bm{v}_{\mathrm{sys}}$ magnitude $v_{\mathrm{sys}}$: when assuming the $\bm{v}_{\mathrm{sys}}$ distribution is isotropic for kinematically old pulsar systems, $v_{\mathrm{sys}}$ is expected to follow a three-component Maxwell distribution (i.e., a weighted sum of three Maxwell distributions) that share the same parameters with the best-fit three-component normal distribution (see Table 8). Namely, under the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption, $v_{\mathrm{sys}}$ should approximately follow the PDF
[…]
@@
where the unit of $p$ and $v_{\mathrm{sys}}$ is $\mathrm{s\penalty 10000\ km^{-1}}$ and $\mathrm{km\penalty 10000\ s^{-1}}$, respectively. In contrast, no analytical solution of the $v_{\mathrm{sys}}$ PDF can be solved from the best-fit Cauchy distribution of $v_{\mathrm{l}}$. Therefore, we adopt Equation 1 as the $v_{\mathrm{sys}}$ PDF in the following analysis, which is illustrated along with its cumulative distribution function (CDF) in Figure 6.
MSP velocities have been previously studied by Hobbs05; Gonzalez11; Matthews16; Shamohammadi24, where their MSP distances were mostly estimated using observed dispersion measures (DMs) and a model (Cordes02; Yao17) of the Galactic free electron distribution, as precise parallaxes were unavailable in most cases. Among these four works, Hobbs05; Gonzalez11 report the mean 3D peculiar velocity of MSPs to be $111\pm 17$$\mathrm{km\penalty 10000\ s^{-1}}$ and $108\pm 15$$\mathrm{km\penalty 10000\ s^{-1}}$, respectively. Both values well agree with the median $v_{\mathrm{sys}}$ (of field MSPs) estimated in this work (i.e., 105 $\mathrm{km\penalty 10000\ s^{-1}}$; see Figure 6). Since no $v_{\mathrm{sys}}$ distribution has been derived in previous studies, a more detailed comparison with our $v_{\mathrm{sys}}$ distribution is not currently feasible.
@@
Regarding the use of Equation 1, we like to clarify two points. Firstly, despite being a three-component Maxwell distribution, Equation 1 is still a uni-modal distribution by definition (see Figure 6). Secondly, though approximating the $v_{\mathrm{sys}}$ PDF by Equation 1, we do not suggest that field MSP systems have multiple formation channels. To further the investigation of formation channels of field MSP systems, population synthesis studies based on various MSP formation channels are needed to generate the theoretical $v_{\mathrm{sys}}$ PDF templates that can be compared with our observation-based counterpart.
@@ 4 Dynamical population synthesis
When deriving the $v_{\mathrm{sys}}$ PDF, we assumed $v_{\mathrm{l}}$ to be statistically stable over time (see Section 3). We tested this “stable-$v_{\mathrm{l}}$” assumption using DPS that simulates the dynamics of field MSP systems in the Galaxy. Specifically, we like to examine if the current (or evolved) $v_{\mathrm{l}}$ distribution stays roughly the same as the initial $v_{\mathrm{l}}$ distribution. We used Equation 1 as the MSP $v_{\mathrm{sys}}$ PDF in the DPS analysis. The MWPotential2014 model, calculated and compiled into the galpy package by Bovy15, was adopted as the Galactic potential used in the DPS. Apart from the $v_{\mathrm{sys}}$ PDF and the Galactic potential, two other ingredients are essential the DPS analysis — the MSP age distribution, and the initial spatial distribution of field MSP systems in the Galaxy. The former ingredient is described in Appendix C.1, while the latter is introduced as follows.
@@ 4.1 The initial Galactic spatial distribution
The initial Galactic distribution of field MSP systems consists of two parts — the birth radial distribution (i.e., the spatial distribution projected to the Galactic plane) and the birth Galactic height distribution (hereafter referred to as the $z_{\mathrm{i}}$ distribution). For the former part, we adopted the pulsar radial distribution established with relatively young pulsars by Xie24, assuming that field MSPs (excluding those originating from globular clusters) are born in pulsar-forming regions similar to those of younger pulsars (such as the spiral arms and the central Galactic region; Xie24).
On the other hand, the $z_{\mathrm{i}}$ distribution of field MSP systems is largely unknown. As we found that the observed $z$ sample (see Table 6) can be well described by a Laplace distribution with the scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc (see Figure 2 and Appendix C.2.1), we assumed that $z_{\mathrm{i}}$ of Galaxy-wide field MSP systems also follow a Laplace distribution with the scale height of $\zeta_{0}$. In DPS, we realized that both the evolved $v_{\mathrm{b}}$ and the evolved $z$ change with $\zeta_{0}$ (see Appendix C.2.2). Therefore, by matching the evolved $v_{\mathrm{b}}$ and the evolved $z$ to the observed counterparts, one can determine $\zeta_{0}$. The DPS procedure that determines the evolved states (including the evolved positions and the evolved velocities) is described in Section 4.2, while the details of the $\zeta_{0}$ determination are provided in Appendix C.2.2. We find that the DPS with $\zeta_{0}=0.32$ kpc renders evolved $v_{\mathrm{b}}$ and evolved $z$ samples that well match the observed counterparts (see Figure 14). Accordingly, we adopt $\zeta_{0}=0.32$ kpc for the Galaxy wide $z_{\mathrm{i}}$ distribution of field MSP systems.
@@ 4.2 The procedure of dynamical population synthesis
Subsequently, we performed orbit integrations with galpy to obtain the evolved positions and velocities after the span of each age in the sample of 13,005 MSP ages described in Appendix C.1. In this way, we compiled $10,000\times 13,005=130$ million evolved states (i.e., positions and velocities for different MSP ages) of simulated MSPs, from which $\sim 4$% evolved states outside the edge of the Galaxy (i.e., beyond 292 kpc from the Galactic centre, Deason20) were discarded. To take into account the selection effect that astrometrically well determined pulsars are predominantly located within $\lesssim 8$ kpc, we removed evolved states that are beyond the 8 kpc distance. Additionally, given that the observed MSPs compiled in this work are $\lesssim 2.5$ kpc from the Galactic plane (see Figure 2), we discarded the evolved states with $|z_{\mathrm{e}}|>2.5$ kpc, where $z_{\mathrm{e}}$ refer to the evolved Galactic heights. The remaining evolved states are $\sim 37$ million in number, which amount to $\sim 30$% of the Galaxy-wide evolved states.
@@ 4.3 Results of dynamical population synthesis
Moreover, from the Galaxy-wide $z_{\mathrm{e}}$ sample (of $\sim 125$ million), we found that the best-fit scale height $\zeta_{1}$ for field MSP systems in the whole Galaxy is 0.68 kpc, which is 31% larger than the observed scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc estimated from relatively nearby MSPs. Interestingly, unlike the observed $z$ sample that is best described by a Laplace distribution (see Appendix C.2.1), the Galaxy-wide $z_{\mathrm{e}}$ sample is best characterized by the Cauchy distribution with 0.48 kpc scale (see Figure 16).
In our analysis, $\zeta_{0}$ and $\zeta_{1}$ were estimated based on the observed $z$ and $v_{\mathrm{b}}$, without accounting for selection effects related to Galactic latitude (also see Appendix C.2.2). As noted in Section 2.1, the observed scale height of binary field MSPs is larger than solitary field MSPs. Therefore, the $\zeta_{0}$ and $\zeta_{1}$ estimated here should reflect the weighted average of the two sub-samples (of binary and solitary field MSPs). In the future, increasing MSP detections across the sky, driven by progressively advanced pulsar search techniques, will enable further refinement of $\zeta_{0}$ and $\zeta_{1}$.